OpenStudy (anonymous):

Rewrite the linear system so that the right terms arranged in columns:ex: 9x-2y+5......2y+-11x+8 Can someone plz help?

OpenStudy (anonymous):

it just says rewrite? is the second one \[2y=11x+8\]?

OpenStudy (anonymous):

yes sorry it is 2y=-11x+8

OpenStudy (anonymous):

what is the first one? \[9x-2y=5\]?

OpenStudy (anonymous):

oh my yes it is so sorry my head is hurting already LOL

OpenStudy (anonymous):

whew ok we have \[9x-2y=5\] \[2y=-11x+8\] is this right?

OpenStudy (anonymous):

if so... rewrite the second one as \[11x+2y=8\] and then put one on top of the other

OpenStudy (anonymous):

to get \[ 9x-2y=5\] \[11x+2y=8\]

OpenStudy (anonymous):

ok ty so much & sorry about the confusion

OpenStudy (anonymous):

then if you feel like solving you can add the two equations together to get 20x=13\] and therefore \[x=\frac{13}{20}\] but i don't know if that is what you are asked to do

OpenStudy (anonymous):

ok your awesome=) thanks again